OpenStudy (love_333):

Fast answer please, medal and fan. In triangle ACE, G is the centroid and BE = 9. Find BG and GE.

OpenStudy (boldjon):

Where is B located? I will need to assume that B is located at the midpoint of side AC, in order to solve the problem. Under this assumption, BE is a median. Along any median, the centroid of a triangle is always located 2/3 of the way from the vertex to the midpoint of the opposite side, so the centroid G is 2/3 of the way from E to B. So we have GE = (2/3)BE. Furthermore, by segment addition, BG + GE = BE. Since BE = 9, GE = (2/3)(9) = 6. Then BG + 6 = 9, which gives BG = 3.